OpenStudy (anonymous):

Expand (x1+2x2+3x3)4

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

Expand (x1+2x2+3x3)4

OpenStudy (anonymous):

to the power of four right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are these subscripts or superscripts?

OpenStudy (anonymous):

\[(x_1+2x_2+3x_3)^4\]or \[(x^1+2x^2+3x^3)^4\]

OpenStudy (anonymous):

subscripts

OpenStudy (mathmale):

hesham, those trying to help you are not picking on you, but rather pointing out that your question is ambiguous. Satellite is trying to figure out what you mean. Could you possibly use the Equation Editor? Lots of people here on OpenStudy could help you learn it.

OpenStudy (anonymous):

\[x_1^4+8 x_2 x_1^3+12 x_3 x_1^3+24 x_2^2 x_1^2+54 x_3^2 x_1^2\]\[+72 x_2 x_3 x_1^2+32 x_2^3 x_1+108 x_3^3 x_1+216 x_2 x_3^2 x_1+144 x_2^2 x_3 x_1+16 x_2^4\]\[+81 x_3^4+216 x_2 x_3^3+216 x_2^2 x_3^2+96 x_2^3 x_3\]

OpenStudy (anonymous):

lol wolfram strikes again! http://www.wolframalpha.com/input/?i=%28x_1%2B2x_2%2B3x_3%29^4

OpenStudy (anonymous):

actually i guess this has to do with elementary symmetric functions, but it will be torture in any case

OpenStudy (mathmale):

If you don't want to use Equation Editor, then please at least indicate exponentiatin like this: x^2 = x squared (a+b)^2 = (a+b) squared

OpenStudy (anonymous):

\[Expand (x1+2x2=3x3)^{4}\]

OpenStudy (anonymous):

answer above

OpenStudy (anonymous):

i want solution be using the multinomial theorem

OpenStudy (anonymous):

step by step plz

OpenStudy (mathmale):

Hesham, as a general rule we do NOT expand an equation such as (a+b=c)^4. Are you positive that there's an " = " sign in there? Are you familiar with "Pascal's Triangle"? I think that 's what you mean by "multinomial theorem." If not, please type out what "multinomial theorem" means, as I'm not familiar with it.

OpenStudy (mathmale):

I'd suggest you do an Internet search for "Pascal's Triangle." That way you'll almost surely find examples of how to use this in expanding powers of binomials.

OpenStudy (anonymous):

you know binomial theorem ?

OpenStudy (anonymous):

binomial theorem for ( a + b)^3

OpenStudy (anonymous):

multinomial theorem for ( a+b +c)^n

OpenStudy (mathmale):

All right; thank you. If you have a "multinomial," such as (a+b+c)^4, you could still find a binomial expansion by grouping any two terms within the parentheses: (a + [b+c])^2 does look like a binomial. Without knowing precisely what you're learning in class and which methods of expansion you'reexpected to use, it's hard to advise you. I still think you'd be well advised to look up and apply Pascal's Triangle for the expansion of powers of binomials.

OpenStudy (anonymous):

ok thanks for you

OpenStudy (mathmale):

Good luck. Hope you'll check out my suggestion that you look up Pascal's Triangle. One of your few remaining options would be to learn the formulas for (a+b)^2 (a+b)^3 (a+b)^4 which are often listed in algebra textbooks.

OpenStudy (anonymous):

(x1 + 2x2 + 3x3) 4 = โˆ‘ n1+n2+n3=4 ( 4 n1 , n2 , n3 ) xn11 (2x2) n2(3x3) n3 The number of terms in the above summation is given by( 4 + 3โˆ’ 1 3โˆ’ 1 ) = ( 6 2 ) = 6 ยท 5 2 = 15 .